OpenStudy (anonymous):

Which property is illustrated by the following statement? Angle X = Angle Y and Angle Z = Angle Y, so Angle X = Angle Z A. Symmetric B. Transitive C. Addition D. Distributive

OpenStudy (anonymous):

This is sort of transitive, but it's not the way they usually ask this type of question. Transitive means that if x = y, and y = z, then x = z Your problem has the second equation reversed, like z = y You can switch it around to read "y = z", but that switch is only possible because of the symmetric property. So the best answer to this question is that both the symmetric property and the transitive property are being used. However, you probably have to guess... if I were guessing, I would guess transitive. Then I would complain about the problem. If you can complain or ask a teacher first, I would recommend doing that, THEN answering the problem given.

OpenStudy (anonymous):

Unfortunately I don't think I can. Seemed strange to me too. It's on an online test I can't back out of until I finish.

OpenStudy (anonymous):

just go with transitive. This is the 2nd "dumb" question I've seen in the last 30 minutes... wish the companies would do a better job of reviewing their tests!!

OpenStudy (anonymous):

(not the 2nd dumb one from you... just saying others are struggling with the wording also)

OpenStudy (anonymous):

I agree! If you mind, can I ask you a few more questions?

OpenStudy (anonymous):

a couple, I have to leave soon. Please post as new questions so others can jump in...

OpenStudy (anonymous):

3x+0=45 3x = 45 What is the reason? A. Addition Property of Equality B. Addition Identity Property C. Multiplication Property of Equality D. Division Property of Equality

OpenStudy (anonymous):

what's your thought?

OpenStudy (anonymous):

Well it's not Multiplication or Division.

OpenStudy (anonymous):

good start. Addition property says you can add the same things to both sides of an equation and it will remain a valid equation. It's like 2 people balanced on a see-saw... hand them each a gallon of milk, and the see-saw stays balanced.

OpenStudy (anonymous):

That would be it then, correct?

OpenStudy (anonymous):

Identity properties are special... they say that adding the identity won't change anything at all.

OpenStudy (anonymous):

Wait.. Now i'm not so sure.

OpenStudy (anonymous):

the 2nd equation says that "adding 0" is like adding "nothing", since 0 is the identity element for addition. Addition Identity says: if y = x then y = x + 0 you can add 0 to one side and it is still OK.

OpenStudy (anonymous):

The multiplication identity is the same idea, but it uses "1" if y = x then y = x * 1 because multiplying by 1 doesn't change anything

OpenStudy (anonymous):

So your answer needs to be "addition identity" because you are dealing with the idea that adding a 0 is like not adding anything at all. Addition property would have been like: 3x + 0 = 45 3x + 0 + y = 45 + y The see-saw was balanced with (3x + 0) on one side, and 45 on the other, and then we added a gallon of milk weighing "y" to both sides, and it was still balanced... that is the "Addition Property"

OpenStudy (anonymous):

Okay, thanks a bunch!